# Convergence of spectral discretization for the flow of diffeomorphisms

### v1

@@ 3 Refined and new regularity estimates
For $k>m+\frac{d}{2}$ it was already shown in [MeP10, Thm. 4.1] by direct estimates that the solution to the geodesic ODE (6) stays in $H^{k-m}(\mathbb{T}^{d};\mathbb{R}^{d})$ if $\rho_{0}$ is.
@@
Theorems 11 and 12 solve a conjecture of [MeP10]: The article considers right-invariant Sobolev metrics (3) of order $m$ on $\mathcal{D}_{\mathrm{id}}^{m+k}$ with $k>0$. Those are known as weak Riemannian metrics since they generate a topology on the tangent spaces $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}=H^{m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ with respect to which the tangent space is not complete. As a consequence, the resulting Riemannian manifold is not metrically complete, but nevertheless the geodesic equation and the Riemannian exponential map may still be well-defined. Indeed, [MeP10, Thm. 4.1] shows that for a smooth metric and $k>m+\frac{d}{2}$ the Riemannian exponential $\exp_{\phi}^{m,k}:T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}\to\mathcal{D}_{\mathrm{id}}^{m+k}$ is a local diffeomorphism, and this is conjectured to be true for any $k\geq 0$. Theorems 11 and 12 confirm this conjecture: Indeed, they show that the restriction of $\exp_{\phi}^{m,0}$ from $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m}$ to $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}$ has range in $\mathcal{D}_{\mathrm{id}}^{m+k}$ (recall that by [BV17] the flow of a $H^{m+k}$-regular velocity lies in $\mathcal{D}_{\mathrm{id}}^{m+k}$), thus $\exp_{\phi}^{m,k}$ is nothing else but this restriction and therefore well-defined. If the metric (3) is three times differentiable, it is even a local diffeomorphism, which follows from the Inverse Function Theorem (noting that differentiability of the metric on $\mathcal{D}_{\mathrm{id}}^{m}$ implies differentiability on $\mathcal{D}_{\mathrm{id}}^{m+k}$).
@@ 4 Spectral space discretization and its properties
In case $f\in H^{k}(\mathbb{T}^{d})$ has no additional regularity one only knows $\|\mathcal{T}_{R}f-f\|_{H^{k}}\to 0$ as $R\to\infty$, but one cannot give a rate. At least one obtains uniform convergence for a continuous family of functions.
@@
where * denotes discrete convolution (to be interpreted in the obvious way if a matrix field is convolved with a vector field), $\hat{D}(\xi)=2\pi i\xi$ is the Fourier multiplier of differentiation, and $\hat{\mathcal{R}}$ and ^=^$\hat{\mathcal{L}}=\hat{\mathcal{R}}^{-1}$ are just the Riesz operator and its inverse expressed as operators on Fourier space. In the following we will assume the Riesz operator and its inverse to be Fourier multipliers so that the relation $v_{t}=\mathcal{R}\rho_{t}$ between velocity and momentum remains after Fourier truncation, $\mathcal{T}_{R}v_{t}=\mathcal{R}\mathcal{T}_{R}\rho_{t}$. Therefore, $\hat{\mathcal{R}}$ and $\hat{\mathcal{L}}$ may simply be interpreted as functions on $\mathbb{Z}^{d}$ in the above.
@@
Since $\hat{P}_{t}$ and $\hat{V}_{t}$ are supported on $Z_{d,R}$, all pointwise multiplications of $\hat{P}_{t}$, $\hat{V}_{t}$, and ^$\widehat{\mathcal{R}}^{R}$ require $O(R^{d})$ floating point operations, likewise the multiplications with $\hat{\mathcal{L}}$ and $\hat{D}$. It remains to compute convolutions of the form $\hat{f}*\hat{g}$ with $\hat{f}$ and $\hat{g}$ supported on $Z_{d,R}$; to this end we extend $\hat{f}$ and $\hat{g}$ to $Z_{d,2R}$ with zeros and simply perform a circular convolution with subsequent truncation to $Z_{d,R}$. Using the Fast Fourier Transform and the Discrete Convolution Theorem this requires $O(R^{d}\log R)$ operations. ∎
@@
As already observed in [ZF19], this discrete approximation may be viewed as geodesic ODE in a finite-dimensional approximate Riemannian Lie group.
@@
The Lie bracket $[v,w]$ satisfies the three characterizing properties of a Lie bracket: bilinearity, antisymmetry, and the Jacobi identity. However, as discussed before, it is not closed in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$, i.e. $v,w\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ does not imply $[v,w]\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$. In contrast, the approximate Lie bracket $[V,W]_{R}$ is closed in $H^{m}_{R}$, but it is not a legitimate Lie bracket as it does not satisfy the Jacobi identity. This is due to a principal obstruction observed by Omori in [Omo78]: There exists no Lie algebra $H^{m}_{R}$ with finite Fourier support and correct Lie bracket. Let us quickly adapt the argument from [Omo78] to our setting, for simplicity restricting to dimension $d=1$: Since the Lie bracket is closed, as soon as the Lie algebra $H^{m}_{R}$ contains trigonometric functions of two different nonzero frequencies, then by repeatedly applying the Lie bracket [,][$\cdot$,$\cdot$][ ⋅ , ⋅ ] one obtains trigonometric functions of arbitrarily high frequencies, which therefore must belong to $H^{m}_{R}$. It does not help to drop the condition on $H^{m}_{R}$ of finite Fourier support, either: There cannot exist any Hilbert Lie algebra containing the constant function and trigonometric functions of arbitrarily high frequencies. Indeed, due to $[[1,\sin(2\pi\xi x)],1]=[2\pi\xi\cos(2\pi\xi x),1]=(2\pi\xi)^{2}\sin(2\pi\xi x)$ the Lie bracket cannot be bounded in any norm on $H^{m}_{R}$. The only exception are the spaces $H^{m}_{R}=\mathrm{span}(1,\sin(2\pi\xi x),\cos(2\pi\xi x))$ with exactly one frequency $\xi\in\mathbb{Z}$; on those, the standard Lie bracket is closed so that $H^{m}_{R}$ is a true Lie algebra.
@@ 5 Convergence of spectral discretization
The first integral is bounded by $\|Bw\|_{H^{0}}^{2}\|\partial_{x_{i}}g\|_{W^{0,\infty}}\lesssim\|w\|_{H^{m}}^{2}\|g\|_{H^{m+1}}$, the second by $\|\partial_{x_{i}}Bw\|_{H^{-1}}\|S\|_{H^{1}}\lesssim\|w\|_{H^{m}}^{2}\|g\|_{H^{m+1}}$ (using [BH21, Thm. 6.1]). Similarly,
@@
of which all summands can be bounded by $\|\rho_{t}\|_{H^{k-m}}\|v_{t}\|_{H^{m}}\|w\|_{H^{m+1-k}}$ (using [BH21, Thm. 6.1] and for $k=0$ proposition 20 in the last summand). ∎
@@
exploiting $[\mathcal{T}_{R}e_{t},\mathcal{T}_{R}e_{t}]=0$ and therefore $[e_{t},\mathcal{T}_{R}e_{t}]=[({\mathcal{I}}-\mathcal{T}_{R})e_{t},\mathcal{T}_{R}e_{t}]=-[({\mathcal{I}}-\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}]$. Using that $H^{m}$ is a Banach algebra as well as lemma 14 and (8) we next estimate
@@
$-\langle[({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}],P_{t}\rangle\leq\|[({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}]\|_{H^{m}}\|P_{t}\|_{H^{-m}}\\ \lesssim\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}\vvvert e_{t}\vvvert_{H^{m}}\vvvert V_{t}\vvvert_{H^{m}}=\vvvert e_{t}\vvvert_{H^{m}}\vvvert V_{0}\vvvert_{H^{m}}\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}\\ \leq\vvvert e_{t}\vvvert_{H^{m}}\vvvert v_{0}\vvvert_{H^{m}}\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}.$
@@
$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$10^{-1}$$10^{0}$$10^{1}$48163264frequency cutoff $R$error $\|e_{1}\|_{H^{2}}$$v_{0}\in H^{3}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{4}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{5}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{6}(\mathbb{T}^{d};\mathbb{R}^{d})$
@@
For a numerical validation of theorem 22 we solve the (space-discrete) geodesic ODE (12) on $[0,1]$ for $R\in\{16,32,64,128\}$ using an explicit $6$-stage Runge–Kutta method (the $5$th order consistent $6$-stage part of the Dormand–Prince method) with step size $2^{-15}$ (by using different step sizes we checked that the error from the time discretization is negligibly small compared to the error from the spectral space discretization). The result for $R=128$ is taken as a substitute for the true solution, and errors for smaller $R$ are computed with respect to that solution. The experiments are performed in $d=2$ space dimensions with $m=3$ and the inner product $(\!(\cdot,\cdot)\!)_{H^{m}}$ defined via the corresponding differential operator $\mathcal{L}=(1-\Delta)^{m}$. To obtain initial conditions $v_{0}$ of different Sobolev regularity we first define a complex-valued $H^{0}$-function $w$ by choosing each Fourier coefficient $\hat{w}(\xi)$ randomly and independently from $[0,1/\sqrt{1+|\xi|^{2}}/\log(2+|\xi|^{2})^{1/2+\epsilon}]$ for $\epsilon=0.1$. Via $\hat{u}(\xi)=\hat{w}(\xi)+\overline{\hat{w}(-\xi)}$ we obtain a real-valued $H^{0}$-function $u$, and we simply set

### v2

@@ 3 Refined and new regularity estimates
For $k>m+\frac{d}{2}$ it was already shown in [MP10, Thm. 4.1] by direct estimates that the solution to the geodesic ODE (6) stays in $H^{k-m}(\mathbb{T}^{d};\mathbb{R}^{d})$ if $\rho_{0}$ is.
@@
Theorems 15 and 16 solve a conjecture of [MP10]: The article considers right-invariant Sobolev metrics (3) of order $m$ on $\mathcal{D}_{\mathrm{id}}^{m+k}$ with $k>0$. Those are known as weak Riemannian metrics since they generate a topology on the tangent spaces $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}=H^{m+k}(\mathbb{T}^{d};\mathbb{R}^{d})$ with respect to which the tangent space is not complete. As a consequence, the resulting Riemannian manifold is not metrically complete, but nevertheless the geodesic equation and the Riemannian exponential map may still be well-defined. Indeed, [MP10, Thm. 4.1] shows that for a smooth metric and $k>m+\frac{d}{2}$ the Riemannian exponential $\exp_{\phi}^{m,k}:T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}\to\mathcal{D}_{\mathrm{id}}^{m+k}$ is a local diffeomorphism [^4] In fact, the proof of the diffeomorphism property in [MP10, Thm. 4.1] does not require smoothness of the metric, but just solvability of the geodesic ODE and differentiable dependence of its solution on the initial data, which holds as soon as the metric is three times differentiable., and this is conjectured to be true for any $k\geq 0$. Theorems 15 and 16 confirm this conjecture: Indeed, they show that the restriction of $\exp_{\phi}^{m,0}$ from $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m}$ to $T_{\phi}\mathcal{D}_{\mathrm{id}}^{m+k}$ has range in $\mathcal{D}_{\mathrm{id}}^{m+k}$ (recall that by [BV17] the flow of a $H^{m+k}$-regular velocity lies in $\mathcal{D}_{\mathrm{id}}^{m+k}$), thus $\exp_{\phi}^{m,k}$ is nothing else but this restriction and therefore well-defined. If the metric (3) is three times differentiable, it is even a local diffeomorphism, which follows from the Inverse Function Theorem (noting that differentiability of the metric on $\mathcal{D}_{\mathrm{id}}^{m}$ implies differentiability on $\mathcal{D}_{\mathrm{id}}^{m+k}$).
Remark 18 (Non-integer exponents).
The results of this section, in particular theorem 15, also hold for non-integer $m$ and $k$ since all employed estimates essentially go back to boundedness of composition with and inversion of Sobolev diffeomorphisms, which can e.g. be found in [IKT13, Lemma B.5-B.6] (on $\mathbb{R}^{d}$, but the extension to bounded domains or manifolds is discussed there, too). Of course, lemma 12 could no longer be applied in theorem 15 to guarantee twice differentiability of the metric, but would have to be replaced by a version in which $B$ is a bounded operator from the fractional Sobolev space $H^{m}$ to $L^{2}$ such that $(w,\phi)\mapsto[B(w\circ\phi^{-1}]\circ\phi$ is $l$ times differentiable (which is the only property used in the proof of lemma 12; [GRR+23, Thm. 1], used in our proof of theorem 15, even just uses differentiability in $\phi$). By [BBC+20, Thm. 4.16 & Thm. 5.1] this is for instance satisfied for any $B\in\mathrm{OPS}_{1,0}^{m}$.
Also the precise control of the constants in lemma 9 and the following carry over: In the proof of lemma 9 one would begin the induction with
$\|f\circ\phi\|_{H^{\theta}}$ $=\int_{\mathbb{T}^{d}}\int_{\mathbb{T}^{d}}\frac{|f(\phi(x))-f(\phi(y))|^{2}}{|x-y|^{d+2\theta}}\,\mathrm{d}x\,\mathrm{d}y$
$=\int_{\mathbb{T}^{d}}\int_{\mathbb{T}^{d}}\frac{|f(x)-f(y)|^{2}}{|\phi^{-1}(x)-\phi^{-1}(y)|^{d+2\theta}}\det D(\phi^{-1})(x)\det D(\phi^{-1})(y)\,\mathrm{d}x\,\mathrm{d}y$
$\leq[(1+C\|\phi-\mathrm{id}\|_{H^{m}})\|f\|_{H^{\theta}}]^{2}$
for $\theta\in(0,1)$ and some $C>0$, exploiting $\det D(\phi^{-1})\leq 1+C\|\phi-\mathrm{id}\|_{H^{m}}$ (cf. the proof of lemma 9) and $|\phi^{-1}(x)-\phi^{-1}(y)|\geq|x-y|/(1+L)$ with $L\lesssim\|\phi-\mathrm{id}\|_{H^{m}}$ the Lipschitz constant of $\phi-\mathrm{id}$ (as can be readily seen after the change of variables $X=\phi^{-1}(x)$, $Y=\phi^{-1}(y)$).
@@ 4 Spectral space discretization and its properties
In case $f\in H^{k}(\mathbb{T}^{d})$ has no additional regularity one only knows $\|\mathcal{T}_{R}f-f\|_{H^{k}}\to 0$ as $R\to\infty$, but one cannot give a rate. At least one obtains uniform convergence for a continuous family of functions (below, recall that $\mathcal{I}$} denotes the identity operator).
@@
where * denotes discrete convolution (to be interpreted in the obvious way if a matrix field is convolved with a vector field), $\hat{D}(\xi)=2\pi i\xi$ is the Fourier multiplier of differentiation, and $\hat{\mathcal{R}}$ and $\hat{\mathcal{L}}=\hat{\mathcal{R}}^{-1}$ are just the Riesz operator and its inverse expressed as operators on Fourier space. In the following we will assume the Riesz operator and its inverse to be Fourier multipliers so that the relation $v_{t}=\mathcal{R}\rho_{t}$ between velocity and momentum remains after Fourier truncation, $\mathcal{T}_{R}v_{t}=\mathcal{R}\mathcal{T}_{R}\rho_{t}$. Therefore, $\hat{\mathcal{R}}$ and $\hat{\mathcal{L}}$ may simply be interpreted as functions on $\mathbb{Z}^{d}$ in the above.
@@
Since $\hat{P}_{t}$ and $\hat{V}_{t}$ are supported on $Z_{d,R}$, all pointwise multiplications of $\hat{P}_{t}$, $\hat{V}_{t}$, and $\widehat{\mathcal{R}}^{R}$ require $O(R^{d})$ floating point operations, likewise the multiplications with $\hat{\mathcal{L}}$ and $\hat{D}$. It remains to compute convolutions of the form $\hat{f}*\hat{g}$ with $\hat{f}$ and $\hat{g}$ supported on $Z_{d,R}$; to this end we extend $\hat{f}$ and $\hat{g}$ to $Z_{d,2R}$ with zeros and simply perform a circular convolution with subsequent truncation to $Z_{d,R}$. Using the Fast Fourier Transform and the Discrete Convolution Theorem this requires $O(R^{d}\log R)$ operations. ∎
@@
As already observed in [ZF19], this discrete approximation may be viewed as geodesic ODE in a finite-dimensional “approximate” Riemannian Lie group. Here and in the following, “approximate” shall only refer to the fact that the associated bracket has a nonzero, but bounded Jacobiator, i.e. the Jacobi identity is only satisfied up to a bounded error. [^5] In the context of deformation theory for Lie groups and Lie algebras it seems that closeness to a Lie algebra can sometimes be quantified by the smallness of the Jacobiator (the deviation from the Jacobi identity), however, in our setting the Jacobiator is not small, but scales like $R^{2}$.
@@
The Lie bracket $[v,w]$ satisfies the three characterizing properties of a Lie bracket: bilinearity, antisymmetry, and the Jacobi identity. However, as discussed before, it is not closed in $H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$, i.e. $v,w\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$ does not imply $v,w]\in H^{m}(\mathbb{T}^{d};\mathbb{R}^{d})$. In contrast, the approximate Lie bracket $V,W]_{R}$ is closed in $H^{m}_{R}$, but it is not a legitimate Lie bracket as it does not satisfy the Jacobi identity. This is due to a principal obstruction observed by Omori in [OMO78]: There exists no Lie algebra $H^{m}_{R}$ with finite Fourier support and correct Lie bracket. Let us quickly adapt the argument from [OMO78] to our setting, for simplicity restricting to dimension $d=1$: Since the Lie bracket is closed, as soon as the Lie algebra $H^{m}_{R}$ contains trigonometric functions of two different nonzero frequencies, then by repeatedly applying the Lie bracket $\cdot$$\cdot$] one obtains trigonometric functions of arbitrarily high frequencies, which therefore must belong to $H^{m}_{R}$. It does not help to drop the condition on $H^{m}_{R}$ of finite Fourier support, either: There cannot exist any Hilbert Lie algebra containing the constant function and trigonometric functions of arbitrarily high frequencies. Indeed, due to $1,\sin(2\pi\xi x)],1]=[2\pi\xi\cos(2\pi\xi x),1]=(2\pi\xi)^{2}\sin(2\pi\xi x)$ the Lie bracket cannot be bounded in any norm on $H^{m}_{R}$. The only exception are the spaces $H^{m}_{R}=\mathrm{span}(1,\sin(2\pi\xi x),\cos(2\pi\xi x))$ with exactly one frequency $\xi\in\mathbb{Z}$; on those, the standard Lie bracket is closed so that $H^{m}_{R}$ is a true Lie algebra.
@@ 5 Convergence of spectral discretization
The first integral is bounded by $\|Bw\|_{H^{0}}^{2}\|\partial_{x_{i}}g\|_{W^{0,\infty}}\lesssim\|w\|_{H^{m}}^{2}\|g\|_{H^{m+1}}$, the second by $\|\partial_{x_{i}}Bw\|_{H^{-1}}\|S\|_{H^{1}}\lesssim\|w\|_{H^{m}}^{2}\|g\|_{H^{m+1}}$ (using lemma 7). Similarly,
@@
of which all summands can be bounded by $\|\rho_{t}\|_{H^{k-m}}\|v_{t}\|_{H^{m}}\|w\|_{H^{m+1-k}}$ (using lemma 7 and for $k=0$ proposition 25 in the last summand). ∎
@@
exploiting $\mathcal{T}_{R}e_{t},\mathcal{T}_{R}e_{t}]=0$ and therefore $e_{t},\mathcal{T}_{R}e_{t}]=[({\mathcal{I}}-\mathcal{T}_{R})e_{t},\mathcal{T}_{R}e_{t}]=-[({\mathcal{I}}-\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}]$. Using that $H^{m}$ is a Banach algebra as well as lemma 19 and (8) we next estimate
@@
$\langle[({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}],P_{t}\rangle\leq\|[({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t},\mathcal{T}_{R}e_{t}]\|_{H^{m}}\|P_{t}\|_{H^{-m}}\\ \lesssim\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}\vvvert e_{t}\vvvert_{H^{m}}\vvvert V_{t}\vvvert_{H^{m}}=\vvvert e_{t}\vvvert_{H^{m}}\vvvert V_{0}\vvvert_{H^{m}}\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}\\ \leq\vvvert e_{t}\vvvert_{H^{m}}\vvvert v_{0}\vvvert_{H^{m}}\|({\mathcal{I}}\!-\!\mathcal{T}_{R})v_{t}\|_{H^{m+1}}.$
@@
$10^{-5}$$10^{-4}$$10^{-3}$$10^{-2}$$^{-1}1001$$^{0}$$10^{1}$48163264frequency cutoff RRerror $\|e_{1}\|_{H^{2}}$$v_{0}\in H^{3}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{4}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{5}(\mathbb{T}^{d};\mathbb{R}^{d})$$v_{0}\in H^{6}(\mathbb{T}^{d};\mathbb{R}^{d})$
@@
Remark 29 (Convergence on $\mathbb{R}^{d}$).
The arguments hold verbatim also when replacing the domain $\mathbb{T}^{d}$ by $\mathbb{R}^{d}$ and interpreting $\mathcal{T}_{R}$ as the truncation operator on continuous Fourier space. However, this is of course of little numerical relevance since the resulting truncated ODE (9) is still infinite-dimensional.
In certain situations the geodesic ODE on $\mathbb{R}^{d}$ can actually be reduced to the geodesic ODE on $\mathbb{T}^{d}$. Indeed, an integral of (6) with momentum $\rho$, velocity $v$ and flow $\phi$ is given by $\rho_{t}=\mathrm{Ad}_{\phi_{t}}^{*}\rho_{0}$, from which one can quickly see that if $\rho_{0}$ has compact support, then so does $\rho_{t}$ for all $t\in[0,1]$ [MTY06, Sec. 4.3.1]. By rescaling and shifting coordinates we may assume $\rho_{t}$ to stay supported within $0,1)^{d}$. Hence, by appropriately choosing the operator $\mathcal{R}_{\mathbb{T}^{d}}$ (the Riesz isomorphism to be used on the torus; the subscript distinguishes it from the original operator $\mathcal{R}$ on $\mathbb{R}^{d}$) one obtains the identical evolution of $\rho_{t}$ on $\mathbb{T}^{d}$ as on $\mathbb{R}^{d}$: Operator $\mathcal{R}_{\mathbb{T}^{d}}$ would be the composition of an extension (of $\rho_{t}$) from $0,1)^{d}$ to $\mathbb{R}^{d}$ by zero (in Fourier space this would mean to interpret the discrete Fourier sum as a sum of Dirac measures and convolving this measure with the Fourier transform $\hat{\chi}$ of the characteristic function $\chi$ of $0,1)^{d}$), an application of $\mathcal{R}$ (which then yields $v_{t}$ on $\mathbb{R}^{d}$), a restriction (of $v_{t}$) to $0,1)^{d}$ (in Fourier space another convolution with $\hat{\chi}$), and a reinterpretation as periodic function (by the Poisson summation formula, the discrete Fourier coefficients would just be the sampling of the continuous ones). Unfortunately, our error analysis only applies in this setting if $\mathcal{R}_{\mathbb{T}^{d}}$ is a Fourier multiplier, which for natural choices (such as $\mathcal{R}=\mathcal{L}^{-1}$ with $\mathcal{L}=(1-\Delta)^{m}$) is not the case.
For initial momentum of unbounded support the situation of course becomes even more involved. For a numerical approximation one would need to periodize momentum or velocity (typically via the Poisson summation formula) in addition to truncating in Fourier space and then analyze the error inflicted by both. A possible direction would be to truncate $\rho_{0}$ in space and investigate the resulting error, after which in a second step the above reduction from $\mathbb{R}^{d}$ to $\mathbb{T}^{d}$ is applied.
For a numerical validation of theorem 27 we solve the (space-discrete) geodesic ODE (12) on $[0,1]$ for $R\in\{16,32,64,128\}$ using an explicit $6$-stage Runge–Kutta method (the $5$th order consistent $6$-stage part of the Dormand–Prince method) with step size $2^{-15}$ (by using different step sizes we checked that the error from the time discretization is negligibly small compared to the error from the spectral space discretization). The result for $R=128$ is taken as a substitute for the true solution, and errors for smaller $R$ are computed with respect to that solution. The experiments are performed in $d=2$ space dimensions with $m=3$ and the inner product $(\!(\cdot,\cdot)\!)_{H^{m}}$ defined via the corresponding differential operator $\mathcal{L}=(1-\Delta)^{m}$. To obtain initial conditions $v_{0}$ of different Sobolev regularity we first define a complex-valued $H^{0}$-function $w$ by choosing each Fourier coefficient $\hat{w}(\xi)$ randomly and independently from $0,1/\sqrt{1+|\xi|^{2}}/\log(2+|\xi|^{2})^{1/2+\epsilon}]$ for $\epsilon=0.1$. Via $\hat{u}(\xi)=\hat{w}(\xi)+\overline{\hat{w}(-\xi)}$ we obtain a real-valued $H^{0}$-function $u$, and we simply set